How many grams of an alloy that is 70% gold should be melted with 20 grams of an alloy that is 10% gold to produce an alloy that is 60% gold ?
Answer by jsmallt9(3759)

For an explanation of where the following equation comes from, click here. Click on the "Back" arrow, "<-" in your browser to return here.

x*0.70 + 20*0.10 = (x+20)*0.60

Now we use algebra to solve it. First we simplify each side:
0.70x + 2 = 0.60x + 12
Subtract 0.60x from each side:
0.10x + 2 = 12
Subtracting 2:
0.10x = 10
Dividing both sides by 0.10:
x = 100

So we mix 100 grams of the 70% gold with the 20 grams of the 10% gold to get 100+20 or 120 grams of 60% gold.